In 2016, U.S. population was 320 million people, and U.S. employment was 160 million people. Using this information and the bar graph above, in 2016, U.S. productivity was approximately _____.

Group of answer choices



$37,500


$50,000


$100,000


$350,000

The GDP for 2016 was $50,000


1
Expert's answer
2021-03-31T07:22:12-0400

Solution:

 

Productivity ="\frac{GDP}{Aggregate \;No. \;of \;labor\; hours}"

GDP = $50,000


Aggregate No. of labor hours ="\frac{U.S. \;employment}{U.S. \; population}"


= "\frac{160,000,000}{320,000,000} = 0.5"

Productivity = "\frac{50,000}{0.5} = 100,000"


U.S. productivity was approximately = $100,000